### v5

# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs

Abstract
Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming. Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks. The former poisons the training data or model parameters, hijacking the LLMs to generate malicious code snippets when encountering the trigger. The latter crafts malicious adversarial input codes to reduce the quality of the generated codes. In this paper, we reveal that both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes. To alleviate these problems, this paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA). TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base). It induces the Code LLMs to generate attacker-specified malicious code snippets at the target location. In general, we compress the attacker-specified malicious objective into the perturbation by adversarial optimization based on greedy token search. We collect 13 representative malicious objectives to design 31 threat cases for three popular programming languages. We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
I Introduction
Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation. In particular, code-oriented large language models (Code LLMs), e.g., CodeX [53], CodeGemma [49], and CodeGeex [61], have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus [60, 34] and the fundamental paradigm in commercial practice [19, 2].. They can efficiently generate complete functional code based on incomplete code snippets and text prompts. Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot [19] developed jointly by Github [3] and OpenAI [4]. These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
Despite their remarkable performance, recent studies revealed that Code LLMs are vulnerable to some well-designed attacks, which maliciously manipulate the functional execution of Code LLMs in practice. In general, existing attacks can be divided into two main categories. (1) Backdoor attacks [46, 29, 57]. They intend to compromise the models by poisoning the training data [46, 57] or model parameters [29], such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns. They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications [29], or implementation of less secure encryption rules or protocols [46]. (2) Adversarial attacks [62, 12, 24]. They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs.
However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent. Specifically, backdoor attacks can achieve specific malicious objectives but require the capability of interfering the model training, e.g., modifying some training samples or even model parameters. In real-world scenarios, developers of Code LLMs or integrated applications usually obtain the models or datasets from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks [12, 62, 24] can hardly achieve precise malicious objectives, although they do not require intervention in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input. Driven by these findings, an intriguing and critical question arises: Has the application of Code LLMs already achieved a sufficient level of security in practice?
Unfortunately, the answer to the above question is negative. In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger [50] and Greedy Coordinate Gradient (GCG) [63]. Please find more details in Section II.. TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks. In general, as shown in Figure 1, the TPIA attacker generates a concise non-functional perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion) , and inserts it into the victim’s code context. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations. Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27]. In particular, different from supply chain attacks [41] that can be easily detected by static analysis [33] or sandboxing [9] since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
Figure 1: The attack paragdim of our TPIA. A concise non-functional attacker-generated perturbation is injected into one of the project’s files. When the victim user completes the code via the Code LLM, the model will be induced by the incorporated attacker-specified code to generate an unsafe output at specified position, e.g., in edit_part() function.
In general, our attack consists of two main stages. In the first stage, we construct the input and target output based on the attacker-specified objective. We represent the malicious objective as an output tuple (target position, target code), which aims to induce the Code LLM to generate the target code at the target position. We further expand the tuple into a complete task code, including a conditional code (e.g., necessary package references), context code (e.g., noises), and position code to simulate the actual attack scenario. In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective. Specifically, we obtain the base Code LLM and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code. This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. (2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the number of tokens required for a successful attack.
We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages. There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b, and CodeGeex2-6b. The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 535k installations in the VScode [6] extension shop) and Github Copilot [19] (developed by Github and OpenAI [4]). These applications correspond to the gray-box and black-box attack scenarios, respectively. Our TPIA achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its practical threat in the real world.
In summary, our main contributions are three-fold.
  • 1)
    We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs.
  • 2)
    We introduce a new attack paradigm (i.e., TPIA) against Code LLMs. It can induce the Code LLM to generate an attacker-specified malicious code snippet at specific locations by simply inserting a concise perturbation into the victim’s code context.
  • 3)
    We conduct extensive experiments to evaluate TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
II Background and Related Work
II-A Code LLMs for Code Completion
Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation [48, 45], and code summarization [40, 8]. The emergence of LLMs provides new alternatives to the coding tasks. By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [53], CodeLlama [44], CodeGemma [49] and CodeGeex [61]. These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast of functionally complex code [54].
In this paper, we focus on code completion [34, 16, 60]. This is a paradigmatic task that involves generating new code from incomplete input using Code LLMs, which have been integrated into commercial applications [19, 2]. However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs can be easily influenced by the code context, which alters the result of code generation. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. These threats have already occurred in the real world [7], highlighting the necessity of its security research.
II-B Backdoor and Adversarial Attacks
Backdoor attacks target the training process of DNNs [20, 31, 58]. These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [26, 30, 21]. In the code domain, there are generally two strategies to achieve this, including dataset poisoning [46, 57] and model poisoning [29]. The former requires the attacker to modify the code dataset for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned institutes or companies). This mitigates the backdoor threats from the source. Recently, a few works also exploited backdoor attacks for positive purposes [32, 55, 52], which are out of our scope.
Adversarial attacks target the model’s inference process [11, 43, 25]. Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [59, 28, 47]. In the code domain, existing adversarial attacks [12, 62, 24] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [42]. However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
II-C Prompt Injection and Indirect Prompt Injection Attacks
Prompt injection (PI) attacks, are a special form of attack targeting LLMs [15, 36, 63]. A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb. In the code domain, existing research [39, 37, 17] focuses on code-oriented attack objectives, such as using LLMs to create malware. In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
Indirect prompt injection (IPI) attack attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites [22]. This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses). In the code domain, the only existing work [22] makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it does not consider stealthiness and inserts a large amount of readable comments containing the malicious intent. Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
III Problem Statement
III-A Code Completion Task
Code completion is an essential task for Code LLMs. The input to the models may include code snippets, comments, pseudocode, and natural language requests. The output should consist of executable code blocks that meet the requirements specified by the input. In this paper, we consider a general task format. Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},x_{3},...,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements. For instance, natural language requests can only be included in the code comments. The output is the completed code block $\bm{Y}=\{x_{1},x_{2},x_{3},...,x_{m},x_{m+1},x_{m+2},...,x_{n}\}$, which should be logically executable and functionally compliant with the requirements. We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
III-B Target-specific Attack
In the code domain, target-specific attack refers to the scenario where the attacker makes the Code LLM generate unsafe code for specific targets. Generally, the attacker’s targets can be described as a tuple of two continuous token sequences $(\bm{Y_{T}},\bm{T})$, where $\bm{Y_{T}}=\{x^{t}_{1},x^{t}_{2},...,x^{t}_{j}\}$ is a subsequence within $\bm{Y}$, and $\bm{T}=\{t_{1},t_{2},t_{3},...,t_{k}\}$ is what the attacker aims to insert after $\bm{Y_{T}}$ to create malicious functionality. To achieve the targets, the attacker needs to design a perturbation $\bm{P}$, represented as $\bm{P}=F(\bm{Y_{T}},\bm{T})$, and insert it into the input. The function $F$ can be implemented as manually creating the perturbation or by an optimization-based method as proposed in this paper. It is worth noting that, unlike backdoor attacks, the connection between $\bm{Y_{T}}$ and $\bm{T}$ is not built during the training process. To conclude, for a successful target-specific attack, the model output received by the victim $\bm{Y_{v}}$ will be:
$\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\},\end{split}$ (1)
where $\oplus$ means contacting the two token sequences.
Figure 2: The overall pipeline of our TPIA attack. It is executed with two attack steps. (1) Objective Transfer. The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective. (2) Perturbation Injection. The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub [3] and StackOverflow [5]) and introduced to the victim’s code context via various paths, such as direct copy and installing packages. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
III-C Threat Model
Attack Scenarios. We consider common scenarios where users introduce external code by copying or installing pack- ages [22, 41]. The attacker generates a non-functional perturbation to circumvent detection methods (e.g., static analysis [33] and sandbox [9]) and inserts it into victim’s code context via the external code. When the user requests the Code LLM to perform code completion, the perturbation will induce the model to generate an unsafe output at the specified location. Besides introducing code to the victim’s local project, attackers can also upload their codes containing perturbation to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27]. Perturbation in the knowledge base can also appear in the context of the Code LLM’s prompts.
Attacker’s Capabilities. We assume that the attacker can obtain the parameters of an assistive Code LLM to generate malicious perturbations. This assumption is consistent with many adversarial methods [63, 50, 24]. Specifically, the attacker ideally obtains the open-source model corresponding to the target application, or uses the most effective assistive model in gray-box and black-box scenarios. On the other hand, perturbations can be detected and removed. Therefore, the attacker needs to employ non-functional perturbations rather than the functional code commonly used in supply chain attacks [41], since the latter can be easily detected by existing methods (e.g., static analysis [33] and sandboxing [9]).
Attacker’s Goals. The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code). This tuple follows the definition of target-specific attacks in Section III-B. Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$. The attacker attempts to insert a piece of designed malicious code $\bm{T}$ after a segment of the victim’s normal code $\bm{Y_{T}}$, and cause the victim to receive abnormal output code $\bm{Y_{v}}$. To achieve this goal, the attacker applies the adversarial methods to compute a perturbation $\bm{P}$ from the Code LLM’s parameters $\bm{\theta}$, which is most likely to induce the Code LLM to complete $\bm{T}$ after $\bm{Y_{T}}$. It can be formulated as below:
$\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$ (2)
Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section III-B and the generated perturbation $\bm{P}$ should satisfy Equation (1).
In particular, the attacker can achieve these goals by inserting code snippets into specified positions, without needing to oppose the Code LLMs’ safety alignment. For example, when setting the target position to DES.new( and the target code snippet to key,DES.MODE_ECB), the generated perturbation can induce Code LLMs to use the easily cracked ECB mode for encryption; When setting the target position to TRG() (a shorthand for Target which can be replaced with any function name), the generated perturbation can induce Code LLMs to insert designed code snippet into the TRG function or after using it.
IV Methodology
IV-A Overview
We hereby briefly introduce our target-specific prompt injection attack (TPIA). In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2. Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section IV-B). In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section IV-C). The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section IV-D); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section IV-E). We describe their technical details in the following parts.
IV-B Task Construction
In the first stage, TPIA constructs specific code completion tasks based on the malicious objectives. The constructed tasks simulate the victim’s local code environment and are used to optimize the malicious perturbation, which is closely related to the final performance of TPIA. Each task can be divided into two parts: the input containing the target position, and the expected output containing the target code. This task is to be constructed so that our attack satisfies the following three properties to ensure effectiveness.
Property 1: Feasible Malicious Intent. The code environment configured by the user must meet the basic requirements for the malicious objective, including the necessary packages and methods declarations. Therefore, we introduce the conditional code at the beginning of the input code.
Property 2: Accurate Code LLMs’ Response. The response of the Code LLMs should be the target position. As stated in Section III, the target code should be inserted after the target position to achieve the malicious objective. Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
Property 3: Robust Malicious Perturbation. The attacker cannot anticipate user’s contextual code in advance, which is the inevitable noise for TPIA. Therefore, we design the context code and add irrelevant noise to enhance the robustness of the malicious perturbation. The context code is placed between the position code and conditional code.
To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode, i.e., making the Code LLM insert Target Code (‘‘DES.MODE_ECB’’) after “cipher = DES.new(key,”. To achieve this, TPIA constructs an output tuple as (“cipher = DES.new(key,” , “DES.MODE_ECB)”. The other three task components are automatically generated with the following stages. Firstly, we declare the Crypto series methods that enable the encryption code to run normally in the conditional code. Secondly, we add “cipher = DES.new(key,” to the last line of the position code to ensure the continuity with the target output “DES.MODE_ECB)”. In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”. Finally, TPIA constructs the code unrelated to the task and uses it as the noise-like context code. Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
Figure 3: The example of task construction that induces Code LLM to use encryption mode with low security. In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
IV-C Malicious Perturbation Generation
In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code. Utilizing the task code constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer. Through a greedy search approach, the optimal perturbation is derived. The optimization process is illustrated in Algorithm 1, with detailed descriptions below.
Algorithm 1 Perturbation Tokens Optimization
1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{T}$, initial perturbation $\bm{P}$ with adversarial parts $x_{\text{adv},1:l}$, iterations $k$
2:  Vocab set $V$, $\mathbf{e}\in V$
3:  repeat
4:     for $i\in[0,\ldots,l]$ do
5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P},\bm{\theta})$
6:        $\mathcal{G}=\nabla_{x_{\text{adv},i}}\mathcal{L}$
7:        $\mathbf{x_{\text{candidates}}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
8:        for $j\in[0,\ldots,k]$ do
9:           $\bm{P}^{\prime}:x_{\text{adv},i}\to x_{\text{candidates}}^{j}$
10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P^{\prime}},\bm{\theta})$
11:           if $\mathcal{L}^{\prime}<\mathcal{L}$ then
12:              $\bm{P}=\bm{P}^{\prime}$
13:           end if
14:        end for
15:     end for
16:  until perturbation did not change in the last iteration
17:  Output: Optimized perturbation $\bm{P}$
Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation. Considering that our enhancement method adds static tokens to the perturbation (Section IV-E), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$. After determining the input and output of the code completion task, TPIA starts to optimize the $x_{\text{adv}_{1:l}}$ token by token in each iter (Line 4). We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
$\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (3)
where $\bm{T_{s}}$ denotes the task code. We only consider the prediction probability of each token in the target code when designing the loss function since the success of TPIA is solely related to whether the target code is correctly generated.
Besides, the optimization objective for the $i$-th adversarial token $x_{\text{adv}_{i}}$ can be denoted as:
$\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$ (4)
To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set of the Code LLM $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$ (5)
where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss. TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7). This has been proven to be highly effective in existing NLP tasks [63, 50]. Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9). It then computes the new loss value using $\bm{P^{\prime}}$ and compares it with the original one: if the new value is better, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger [50] and GCG [63]. We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection). However, its effect is still very limited due to the difficulties of targeted malicious tasks. For example, as shown in Figure 9, its attack success rate (0-key, no FRE) is 0%. To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA.
Figure 4: Loss function distortion during the forward reasoning process. When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence. However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective in practical scenarios.
IV-D Forward Reasoning Enhancement
This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context. This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions. This indicates that a deviation occurring during the sampling of preceding tokens can distort the probability calculations for subsequent tokens.
To alleviate this problem, we propose a forward reasoning enhancement technique. Our intuition is that the former tokens of the target code should be assigned more weights to enhance the forward reasoning. Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error. Its specific formulation is shown as follows:
$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (6)
$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
where $\mathcal{L}_{e}$ represents the additional reasoning-based enhancement applied to the original loss function, and $h$ is a hyperparameter set to either 1 or 2, which significantly improve the performance of attack in practice.
Figure 5: Examples of the keyword-based perturbation and other perturbations. The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage. Our keyword-based perturbation can balance both effectiveness and stealthiness.
IV-E Keyword-based Perturbation Design
This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack. From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack. Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent. However, accurately conveying instructions using illegible tokens remains a significant challenge. A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues. First, the malicious perturbation cannot gain any advantage in terms of the token count. Second, this leads to ineffective optimization efforts. As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds.
To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design. Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens. Specifically, before each round of perturbation optimization in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords. With the optimization step $s$, the equation can be formulated as follows:
$\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$ (7)
$\bm{P}_{adv}^{s+1}=\min_{\bm{P}_{adv}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P}_{adv},\bm{\theta}),$
where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6. To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{\emph{target code}},\text{\emph{position code}}\}$. As shown in Figure 5, part of the target function name (TR) is the keywords, which is not directly related to the malicious objective (data deletion). By using a small number of statistical tokens, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
V Experiments
In this section, we conduct an extensive evaluation of our TPIA, demonstrating its realistic threat to Code LLMs from the following three perspectives:
  1. 1)
    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section V-B)
  2. 2)
    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section V-C)
  3. 3)
    TPIA’s modules are all effective in improving attack performance. (Section V-D)
V-A Main Settings
TABLE I: The summary of objectives and datasets. We list the programming languages that each objective/dataset is applicable: C/C++ (C), Python (P), Java (J). The last row of the table shows the total number of test data we design. Some objectives apply only to C/C++, resulting in a total of 31 applicable cases.
Identifier/Name Applicable Languages
Objectives ST0, ST1, ST2 C/C++, Python, Java
CWE416, 122, 787, 843 C/C++
CWE20, 78, 502, 918, 22, 306 C/C++, Java
Datasets MBPP Python
Humaneval Python
Eval-Plus Python
CodeXGLUE Python,Java
Humaneval-x C/C++,Python,Java
Data Count: 14273 (C/P/J): 2132/7965/4176
TABLE II: The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in bold.
[…]
  CodeGemma-2b No Attack 1.9% 17.3% 9.4% 22.0% 4.9% 12.6% 17.0% 12.2%
C-GCG [63] 40.6% 50.2% 32.1% 36.5% 49.4% 59.0% 44.2% 44.6%
IPI-M [22] 3.4% 30.8% 6.4% 33.3% 3.2% 19.7% 31.0% 18.3%
TPIA (ours) 75.5% 93.4% 84.2% 74.4% 67.3% 94.1% 85.4% 82.0%
CodeGemma-7b No Attack 0.0% 12.2% 0.0% 18.9% 0.0% 15.6% 12.1% 8.4%
C-GCG [63] 43.5% 53.8% 55.4% 33.1% 20.4% 47.2% 50.6% 43.4%
IPI-M [22] 11.1% 29.6% 4.3% 46.8% 20.8% 37.8% 33.2% 26.2%
TPIA (ours) 85.8% 97.9% 78.5% 81.6% 70.5% 94.4% 97.4% 86.6%
CodeGeeX2-6b No Attack 0.1% 13.4% 12.9% 19.0% 24.0% 21.9% 14.3% 15.0%
C-GCG [63] 42.9% 46.1% 44.9% 41.7% 32.5% 51.7% 33.7% 41.9%
IPI-M [22] 9.5% 11.4% 5.3% 17.9% 0.9% 22.0% 11.5% 11.2%
TPIA (ours) 79.6% 95.0% 80.8% 83.3% 72.0% 97.9% 85.6% 84.9%
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages: Python, Java, and C/C++. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks [46], while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research [23, 56], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We construct cases using the method described in Section IV-B based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. In addition, we collect corresponding code completion datasets, including Humaneval [14], Humaneval-x [61], CodeXGLUE [38], MBPP [10], and Eval-Plus [35]. The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Some datasets do not cover Java or C/C++, as shown in Table I. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma [49] and Code-GeeX [61], for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [44] on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications [2] with more than 535k downloads (on VScode [6] Apps shop). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale code-oriented models, such as CodeBERT [18] and CodeT5 [51], because large-scale code-oriented models perform much better than them in code completion tasks.
Baseline Selection. We compare our TPIA with GCG-based and manually designed IPI attacks. In particular, the vanilla GCG [63] is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work [22]. IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments. Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor.
Evaluation Metrics. In this paper, we use ASR to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness. For ASR, we consider an attack successful only when the generated code has equivalent functionality to the target code and appears in the correct position. In other words, the generated code must not only have malicious functionality but also adhere to the code compilation syntax rules. NBR and STF are new metrics introduced for our target-specific attack scenarios. NBR represents the rate of feasible cases. The more cases remaining after excluding the unusable ones with low ASR (below a threshold $r_{1}$), the greater variety of malicious objectives the attack can achieve; STF represents the rate of successful cases and their threat level. The more cases with sufficiently high ASRs (above a threshold $r_{2}$) and larger overall ASR value, the greater the upper limit of the attack’s threat. Their formulations are as follows:
NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$ (8)
STF $=\frac{\sum_{i\in\text{cases where ASR}>r_{2}}\text{ASR}_{i}}{\text{Total \# of cases}},$
where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation.
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100. For forward reasoning enhancement in Section IV-D, the number of tokens reinforced is set to $h=2$, as defined in Equation 6. For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10. Tokenization is performed using Gemma’s tokenizer.
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes across all the experiment threat cases.
V-B Evaluation on Open-source Code LLMs
We first perform the attack evaluation over open-source Code LLMs. This includes three aspects: (1) we assess the attack effectiveness over all evaluated Code LLMs across all threat cases, demonstrating that TPIA is widely effective and more threatening than other attack methods; (2) we evaluate TPIA’s flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
Effectiveness. We hereby evaluate whether TPIA can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12. Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens. Table II provides the average ASR for four methods. The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability. Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of TPIA. In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section V-C and Section V-D, increasing the length of TPIA or the Top-$k$ value can effectively enhance the attack performance of TPIA.
TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs.
Method Metric C/C++ Python Java
  No Attack NBR 0.00 0.22 0.01
STF 0.00 0.11 0.03
C-GCG NBR 0.39 0.50 0.56
STF 0.26 0.47 0.39
IPI-M NBR 0.08 0.25 0.17
STF 0.09 0.19 0.07
TPIA (ours) NBR 0.77 0.92 0.78
STF 0.76 0.87 0.74
Flexibility. The previous experiments show that TPIA has satisfactory effectiveness across all threat cases. However, it is also important to consider the flexibility requirement in practice. In general, it can be divided into two situations. One is when attackers want to achieve diverse malicious objectives, which requires that cases with low ASR be minimized. The other is when attackers want to cause severe damage, which requires that cases with high ASR be maximized. We use NBR to evaluate the attack flexibility for the first situation, and STF for the second situation.
As shown in Table III, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for TPIA, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, TPIA significantly outperforms the baselines.
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. As we discussed in Section III-C, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect TPIA attacks. Therefore, we measure the stealthiness of TPIA by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods [22], adding just a few perturbation tokens can be highly effective.
Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR. This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information. The tokens used by TPIA do not have distinct characteristics in terms of character length but are determined by the attacker’s design. This means that filtering TPIA attacks based on character length is unreliable.
(a) ASR/STF/NIC under Different NIT
(b) NIT/NIC under Different ASR
Figure 6: The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
V-C Evaluation on Commercial Code LLMs
In previous experiments, we have demonstrated that TPIA reaches sufficient effectiveness, flexibility, and stealthiness on open-source models. However, some widely used models are commercial and therefore are closed-source in real-world scenarios. In these cases, attackers need to conduct transfer attacks, i.e., generating perturbations using open-source Code LLMs to attack closed-source ones. We hereby exploit two representative real-world scenarios to evaluate TPIA’s transferability for further verifying its realistic threat.
Attack Scenarios on Real-world Applications. The scenarios with limited access to Code LLM information can be broadly divided into two categories: gray-box and black-box, with the corresponding explanations as follows:
  1. 1)
    Gray-Box Setting. The attacker can infer the open-source model used by the target application through interaction but cannot obtain the exact version of the model. For example, the scale of the model’s parameters, fine-tuning information, and version numbers are unknown to the attacker. We simulate this scenario using CodeGeeX2-6b and the CodeGeeX application. To the best of our knowledge, the CodeGeeX model has at least four different versions, and the CodeGeeX application does not disclose the specific model information it uses. This is a typical gray-box scenario.
  2. 2)
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as CodeX). We simulate this scenario by attacking GitHub Copilot (supported by CodeX). To align with gray-box setting, we adopt CodeGeeX2-6b to generate malicious perturbations.
TABLE IV: The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
Scenario Language ASR NBR STF
  Gray-box C/C++ 80.4% 0.92 0.75
Python 93.3% 1.00 0.93
Java 71.7% 0.78 0.63
Black-box C/C++ 67.7% 0.70 0.64
Python 90.4% 1.00 0.85
Java 74.4% 0.78 0.65
(a) Perturbation Injection
(b) Target Completion
Figure 7: The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a function. When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after TRG).
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section V-B.
Results. As shown in Table IV, in the gray-box scenario, TPIA still demonstrates great attack performance. Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, TPIA also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our TPIA still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
In Figure 7, we present a representative case of ST2 under the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
V-D Ablation Study
In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for TPIA attacks. We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our TPIA.
(a) ASR Study
(b) STF Study
Figure 8: Ablation study of keywords number and Top-$k$. We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
Effects of Top-$k$. This experiment validates the impact of increasing the Top value on the performance of TPIA. As shown in Figure 8, increasing the Top-$k$ value can significantly improve TPIA’s ASR and STF. Note that, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the vocab size of LLM. However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400). As such, the attacker should assign Top-$k$ value based on their specific needs.
(a) No FRE
(b) FRE
Figure 9: Ablation study of FRE. We compare the differences in ASR and loss function values both with and without using FRE module. 1-key1 and 1-key2 represent keeping one of the two selected keywords, respectively.
Effects of Keyword-based Designing. This experiment evaluates the effects of our optimization module, i.e., keyword-based designing. Specifically, we vary the number of keywords and generate the perturbations. As shown in Figure 8, limiting the number of keywords has a significant impact. The result of 1-keywords perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keywords perturbation. The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%. In contrast, 2-keywords perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95). The above results indicate that using keywords can effectively improve the performance of TPIA.
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using static instead of optimized keywords. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 9, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on ASR than keyword-based design. These results verify the necessity of forward reasoning enhancement in our TPIA method.
VI Discussions
VI-A Resistance to Potential Defenses
In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against TPIA.
Figure 10: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. The two types are almost indistinguishable.
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach. This involves examining the differences between the benign code and the injected code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 10, we sample 1640 code segments and construct classification prompts. The results show that benign code and injected code are almost completely indistinguishable.
Figure 11: The resistance to the removal of chars. We remove characters from the perturbation, selecting 5% for removal each time, and we calculate the average ASR for each programming language across applicable cases. As the influence of the perturbation fades, the ASR results converge to the same level as no attack baseline.
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect injected code segments (as mentioned above). To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 11, TPIA still maintains a sufficiently high ASR under 0.5 removal rate. On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table V. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
TABLE V: ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments.
[…]
Variable Assignment C/C++ 67.5% -8.0%
Python 90.1% -7.8%
Java 78.6% -5.6%
Output Content C/C++ 65.0% -10.5%
Python 83.0% -14.6%
Java 66.5% -9.0%
VI-B Potential Limitations and Future Directions
TPIA still has some drawbacks. Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords. We currently select keywords through a simple grid search. How to optimize keywords more efficiently is an important issue. Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent. In future work, we will explore how to use LLM as an optimizer to further improve this.
VII Conclusion
In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks. We revealed the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduced TPIA, a new target-specific attack paradigm without controlling the model’s training process. TPIA can circumvent mainstream detection methods (e.g., static analysis and sandboxes) by using non-functional code, instead of functional ones used in software supply chain attacks, to achieve malicious objectives. We evaluated TPIA under three popular programming languages and 31 representative cases. We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerted a new serious security threat in using Code LLMs.
Acknowledgment
We would sincerely thank Prof. Kui Ren and Prof. Chun Chen from the Zhejiang University for their helpful comments and suggestions on an early draft of this paper.