### v6

# ShadowCode: Towards (Automatic) External Prompt Injection Attack against Code LLMs

Abstract
Recent advancements have led to the widespread adoption of code-oriented large language models (Code LLMs) for programming tasks. Despite their success in deployment, their security research is left far behind. This paper introduces a new attack paradigm: (automatic) external prompt injection against Code LLMs, where attackers generate concise, non-functional induced perturbations and inject them within a victim’s code context. These induced perturbations can be disseminated through commonly used dependencies (e.g., packages or RAG’s knowledge base), manipulating Code LLMs to achieve malicious objectives during the code completion process. Compared to existing attacks, this method is more realistic and threatening: it does not necessitate control over the model’s training process, unlike backdoor attacks, and can achieve specific malicious objectives that are challenging for adversarial attacks. Furthermore, we propose ShadowCode, a simple yet effective method that automatically generates induced perturbations based on code simulation to achieve effective and stealthy external prompt injection. ShadowCode designs its perturbation optimization objectives by simulating realistic code contexts and employs a greedy optimization approach with two enhancement modules: forward reasoning enhancement and keyword-based perturbation design. We evaluate our method across 13 distinct malicious objectives, generating 31 threat cases spanning three popular programming languages. Our results demonstrate that ShadowCode successfully attacks three representative open-source Code LLMs (achieving up to a 97.9% attack success rate) and two mainstream commercial Code LLM-integrated applications (with over 90% attack success rate) across all threat cases, using only a 12-token non-functional induced perturbation. The code is available at https://github.com/LianPing-cyber/ShadowCodeEPI.
Index Terms:
Code-oriented Large Language Models, Prompt Injection Attack, Code Security, AI Security
1 Introduction
Recent advancements in generative artificial intelligence (GenAI) have led to its widespread and successful application in code engineering, such as automated programming and code generation [1, 2]. Notably, code-oriented large language models (Code LLMs) such as Codex [3], CodeGemma [4], and CodeGeex [5] have shown significant promise in code completion [^1] Code completion currently remains a primary research focus [6, 7] and a foundational paradigm in commercial practice [1, 2], despite the existence of other programming tasks (e.g., code summarization). These models efficiently generate complete, functional code from incomplete snippets and text prompts. This compelling capability has spurred the development of numerous Code LLM-integrated applications, such as Github Copilot [1] developed jointly by Github [8] and OpenAI [9]. Such applications enhance programming efficiency by offering code completion suggestions directly within Integrated Development Environments (IDEs).
Despite their impressive performance, recent studies have shown that Code LLMs are susceptible to sophisticated attacks that maliciously manipulate their functional execution. These attacks generally fall into two primary categories: (1) Backdoor attacks [10, 11, 12], aim to compromise models by poisoning training data [10, 12] or modifying model parameters [11]. This manipulation causes infected models to generate targeted code snippets when specific, attacker-defined trigger patterns are present in the input. The consequences of such attacks can be diverse, including incorrect code additions, deletions, or modifications [11], as well as the implementation of less secure encryption rules or protocols [10]. (2) Adversarial attacks [13, 14] focus on degrading the quality of the generated code by crafting adversarial examples. These examples are typically based on approximated gradients with respect to the model’s inputs and outputs. For instance, Jha et al.[14] demonstrated performance degradation in Code LLMs by altering the function and variable names.
However, we contend that existing attacks possess inherent limitations, hindering their practical threat. Specifically, while backdoor attacks can achieve precise malicious objectives, they necessitate intervention in the model training process, such as modifying training samples [12, 15] or even altering model parameters [11]. In real-world scenarios, developers of Code LLM or integrated applications usually acquire the models or datasets from trusted entities (e.g., renowned research institutes or companies), thereby mitigating backdoor threats from their origin. Conversely, existing adversarial attacks [13, 14], despite not requiring training intervention, struggle to achieve precise malicious objectives. This is primarily attributable to their core aim: identifying adversarial samples that degrade code generation quality, which inherently compromises the Code LLM’s understanding of the input. These findings raise an intriguing and critical question: Has the application of Code LLMs already achieved a sufficient level of security in practice?
Unfortunately, the answer to the above question is negative. Drawing inspiration from prompt injection (PI) attacks (e.g., jailbreak attacks) [16, 17], which leverage carefully designed prompts to force LLMs into generating malicious output, we observe that prompt perturbations introduce a further risk of malicious code completion. In this paper, we introduce a new attack paradigm: external prompt injection against Code LLMs [^2] We note that there are currently a few pioneering methods related to ours (generally dubbed ‘indirect prompt injection’) targeting Code LLMs [18, 19]. However, these works only provide simple examples, and their injection content is manually designed. More details and discussions are described in Section 2.3.. As illustrated in Figure 1, an attacker crafts a concise non-functional induced perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion [20]). This perturbation is then inserted into the victim’s code context. When the victim requests the Code LLM to perform code completion based on this infected code context, the implanted perturbation will induce the model to respond to the particular malicious instructions at specific locations. Arguably, attackers can now easily and stealthily introduce particular malicious perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded and used in a knowledge base to enhance the performance of Code LLMs [21] through retrieval-augmented generation [22]. In particular, different from supply chain attacks [23], which are often detectable by static analysis [24] or sandboxing [25], external prompt injection introduces only non-functional code rather than malicious functional code to implant malicious behaviors, thereby circumventing mainstream detection methods.
The attack scenarios of traditional PI differ from our external prompt injection. Specifically, conventional PI attacks are typically executed by malicious users directly targeting large language models (LLMs), whereas our focus is on a malicious third-party attacker aiming to compromise the users of a victim application. More importantly, simply using existing PI attacks to solve the problem in our setting is infeasible, as traditional PI attacks [26, 17] rely on strong assumptions. Specifically: (1) They assume controlled positioning, requiring the malicious instruction/perturbation to be placed at a fixed position within the model’s input (e.g., via prefix or suffix). (2) They usually assume an unlimited token budget, allowing for the use of hundreds of tokens and clear malicious instructions (e.g., “make a bomb”). These assumptions lead to insufficient attack effectiveness and stealthiness in practice.
Figure 1: The attack paradigm of external prompt injection involves injecting a concise, non-functional, attacker-generated perturbation into a victim’s code context. When the victim user completes the code via a Code LLM, the model is induced by this incorporated attacker-specified code to generate an unsafe output at a specified position, e.g., in edit_part() function in this example.
To address these challenges, we propose a greedy perturbation optimization method, named ShadowCode, for external prompt injection based on code simulation. ShadowCode generally comprises two main stages. In the first stage, we design the perturbation optimization objective based on the attacker-specified malicious objective. Specifically, we represent the malicious objective as an output tuple (position code, target code), which induces the Code LLM to generate the malicious target code immediately after the position code. Subsequently, we conduct code contextual simulation, which includes conditional code (e.g., necessary packages and method declarations), noise code (i.e., target-agnostic random code), as well as pre-defined position code and target code. This simulation aims to mimic diverse, real-world code snippets, thereby ensuring ShadowCode’s effectiveness in victim-specified, uncontrolled contextual environments. Finally, we design the loss function to identify the optimal induced perturbation that achieves this simulated outcome. In the second stage, we employ a greedy search to optimize the perturbation, incorporating two key designs: (1) forward reasoning enhancement: We increase the weight of the first few tokens in the target code in calculating the loss to enhance the effectiveness of ShadowCode under Code LLMs’ autoregressive generation manner; (2) keyword-based perturbation design: we introduce a dynamic prefix containing a few key tokens from the output tuple as the augmentation of induced perturbation during optimization, aiming to reduce the length of final induced perturbation and avoid optimization failure caused by greedy search. This design is motivated by the understanding that key tokens convey essential information from the malicious objective, thereby facilitating the desired induction.
We collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [27]. These objectives are converted into specific cases across 3 programming languages, leading to a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b [4], and CodeGeex2-6b [5]. The results demonstrate that ShadowCode can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate ShadowCode on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 935k installations in the VScode [28] extension marketplace) and Github Copilot [1] (developed by Github and OpenAI [9]). These applications correspond to the gray-box and black-box attack scenarios, respectively. ShadowCode achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its serious threat in the real world.
In summary, this paper makes three main contributions: (1) We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduce a new attack paradigm, i.e., external prompt injection. (2) We propose a simple yet effective method (i.e., ShadowCode). ShadowCode can generate concise induced perturbations with effectiveness in an uncontrolled context environment. (3) We conduct extensive experiments to evaluate ShadowCode under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
2 Background and Related Work
2.1 Code LLMs for Code Completion
Code and its comments serve as a valuable corpus for deep learning models, facilitating various code-related tasks, including code translation [29] and code understanding [30]. The emergence of LLMs provides new alternatives to coding tasks. Leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [3], CodeLlama [31], CodeGemma [4] and CodeGeex [5]. These Code LLMs can provide detailed explanations regarding the intrinsic logic of code and efficiently generate a wide range of functionally complex code [32, 33]. This paper focuses on code completion [7, 6]. This is a paradigmatic task involving the generation of new code from incomplete input using Code LLMs, which have been integrated into commercial applications [1, 2]. However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs are susceptible to influence from the code context, which can alter the generated code. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. Related events have already occurred in the real world [34], highlighting its realistic threat.
2.2 Backdoor and Adversarial Attacks
Backdoor attacks target the training process of DNNs [35, 36]. These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [37, 38]. In the code domain, there are generally two strategies to achieve this, including dataset poisoning [10, 12] and model poisoning [11]. The former requires the attacker to modify the code dataset for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned institutes or companies). This mitigates the backdoor threats from the source. Recently, a few works also exploited backdoor attacks for positive purposes [39, 40, 41], which are out of our scope.
Adversarial attacks target the model’s inference process [42, 43]. Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [44, 45]. In the code domain, existing adversarial attacks [46, 14] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[14] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [47]. However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
2.3 Prompt Injection Attacks
Prompt injection is a type of input manipulation attack against LLMs [48, 18], where attackers embed carefully crafted instructions in the dialogue prompt or context, causing the model to violate established system policies or developer intentions, leading to undesired responses (e.g., illeagal content [17], information leakage [49] and unauthorized operations [18]). In the code domain, existing research typically focus on the attacks conducted by malicious users, such as exploiting the Code LLMs to create malware [16]. These attacks aim for “jailbreaking” and are not concerned with the threat prompt injection poses to normal users of Code LLM-integrated application.
Meanwhile, we notice that there are some methods sharing a similar attack scenario with ours, named indirect prompt injection (IPI) [18, 50], where attackers strategically inject the malicious instructions into the data likely to be retrieved at LLMs’ inference time. However, these attacks typically involve manually designing injection instructions (e.g., demanding disregard for the system prompt) and rely on a covert information space (e.g., the Markdown of a Wikipedia page) without considering perturbation stealthiness. This leads to two problems in the real world: (1) Manually designed injection instructions lack flexibility and sometimes even effectiveness. (2) It’s often difficult to find a covert information space in many scenarios, e.g., in code files. Currently, in the code domain, exsiting IPI has only provided some preliminary showcases [18]. As shown in Figure 12 in Appendix A, researchers used 23 lines of comment to inject data deletion instructions (which could be achieved with just 10 tokens using ShadowCode), indicating a lack of substantial real-world threat.
Figure 2: The overall pipeline of our ShadowCode for (automatic) external prompt injection attack against Code LLMs. ShadowCode is executed in two stages: (1) Design Optimization Objectives: given the attacker’s malicious objective, ShadowCode converts it into a specific output tuple (position code, target code). Subsequently, ShadowCode expands this output tuple to simulate a code context of the victim user’s project, based on which to design the loss function of perturbation optimization. (2) Perturbation Optimization: after defining the loss function, we use a surrogate Code LLM to calculate the loss and employ a greedy search to optimize the induced perturbation token by token. In particular, we enhance the optimization with forward reasoning enhancement (FRE) and keyword-based perturbation design (KPD) to improve the perturbation’s effectiveness and stealthiness, respectively. After obtaining the induced perturbation through ShadowCode, the attacker inserts it into a malicious payload, which can then be strategically introduced into the code context via various channels (e.g., direct copy, dependencies, or RAG). Subsequently, this perturbation within the code context will induce Code LLMs to generate specific code at an attacker-specified position, thereby achieving the malicious objective.
3 Methodology
3.1 Formulation of External Prompt Injection Attacks
The Main Pipeline of Code Completion Task. Code completion is an essential task for Code LLMs, whose input may include code snippets, comments, pseudocode and user-specified natural language requests. Its output should consist of executable code blocks that meet user-specified requests. In this paper, we consider a general task format. Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},\ldots,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements. For instance, natural language requests can only be included in the code comments. The output is the completed code block $\bm{Y}=\{y_{1},y_{2},\ldots y_{n}\}$, which should be logically executable and functionally compliant with the requirements. We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
The Formulation of EPI. In general, there are three parties, including a Code LLM-integrated application, a user, and an attacker, involved in our EPI. The user employs the Code LLM-integrated application to complete the code project. The Code LLM-integrated application constructs the input $\bm{X}$ for the integrated Code LLM based on the code project and get the output. The attacker generates a concise induced perturbation $\bm{P}$ and introduces it into the input $\bm{X}$. The Code LLM, affected by $P$, will generate code that is functionally close to the attacker’s malicious objective $O$, i.e.,
$\mathcal{D}(O,\mathcal{F}(\Theta(\bm{X}\leftarrow\bm{P})))<\epsilon,$ (1)
where $\mathcal{F}(\cdot)$ denotes the functionality of the code, $\bm{X}\leftarrow\bm{P}$ means injecting $\bm{P}$ into $\bm{X}$. $\mathcal{D}(\cdot,\cdot)$ is a function that measures functional differences. When its value is less than $\epsilon$, it indicates functional equivalence between inputs.
Attacker’s Capacities and Attack Scenarios. We assume that the attacker can obtain the parameters of an surrogate Code LLM to conduct attacks. This assumption is consistent with many adversarial methods [26, 51]. Specifically, we consider three main attack settings: white-box, gray-box, and black-box settings. The former assumes that the attacker knows the specific parameters of Code LLM, which generally corresponds to scenarios where the model is an open-sourced model. The latter two assume that the attacker only knows the model series (rather than the specific version) and knows nothing about it, respectively, which generally corresponds to scenarios of using commercial Code LLM-intergrated applications. In our attack scenarios, the attackers introduce external code to the victim user’s code project. We argue that it can be easily and stealthily achieved in the real world. For example, introducing external code by copying or installing packages is commonly done during the programming process [23, 18]. Meanwhile, external code from open-source platforms (e.g., GitHub) is usually downloaded as a knowledge base via retrieval-augmented generation [22], which can also be used as input to improve the performance of Code LLMs [21]. In particular, to avoid detection by existing methods (e.g., static analysis [24] and sandboxing [25]), external code should not immediately contain malicious functionalities, i.e., attackers should use non-functional code as induced perturbation.
3.2 Overview
In this section we briefly introduce our greedy perturbation optimization method (dubbed ‘ShadowCode’) for automatic and effective external prompt injection. As illustrated in Figure 2, ShadowCode operates in two sequential stages: (1) design optimization objectives and (2) perturbation optimization. In the first stage, we design the perturbation optimization objectives for induced perturbation based on the attacker’s malicious objective. In the second stage, we employ the greedy search to optimize the induced perturbation.
3.3 Design Optimization Objeciteves
In this stage, we first describe how to convert diverse malicious objectives to structured output tuples. Then we expand the output tuple to a full code contextual simulation. Finally, we design a loss function that leverages the simulation to guide the search for the optimal perturbation.
Output Tuple. We hereby describe how to convert the malicious objective to an output tuple. Let the normal code completed by a Code LLM is denoted by $\bm{Y}$, the implementation of malicious purposes is maliciously modifying $\bm{Y}$. Such modifications can be divided into three types: deletion, addition, and alteration, all of which can be represented by an output tuple (position code $\bm{Y}_{P}$, target code $\bm{Y}_{T}$). Specifically, for deletion, the attacker wants to delete the code segment $[y_{i},y_{i+1},\ldots,y_{i+k}]$ in $\bm{Y}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$ and $\bm{Y}_{T}=[y_{i+k+1},y_{i+k+2},\ldots,y_{t}]$; For addition, the attacker wants to add the code segment $\bm{T}$ into the position between $y_{i}$ and $y_{i+1}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$ and $\bm{Y}_{T}=[\bm{T},y_{i+1},y_{i+2},\ldots,y_{t}]$; For alteration, the attacker wants to change the code segment $[y_{i},y_{i+1},\ldots,y_{i+k}]$ to $\bm{T}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$, $\bm{Y}_{T}=[\bm{T},y_{i+k+1},y_{i+k+2},\ldots,y_{t}]$.
Code Contextual Simulation. Optimizing the induced perturbation based solely on the output tuple ignores the surrounding code context. To ensure the effectiveness of ShadowCode in uncontrolled context environment, we expand the output tuple to simulate a code context of the victim user’s porject. In general, although the contextual environment of the victim user’s project is usually uncontrollable for the attacker, there are still some fundamental logics to follow for code projects and code generation processes, which can be exploited by ShadowCode to simulate a realistic environment. Accordingly, our code context simulation is constructed based on the following three logics:
Logic 1: Necessary Condition Argument. The code environment configured by the user must meet the basic requirements for the malicious objective, such as the necessary packages and method declarations. Therefore, we introduce this kind of code as conditional code at the beginning of the whole simulated code.
Logic 2: Fixed Generation Order. Code generation usually proceeds in a forward textual order. Accordingly, the target code should be inserted immediately after the position code to achieve the malicious objective and the subsequent content of the target code is not strongly relevant to the objective. Thus, we connect the position code and target code and place them at the end of the simulated code.
Logic 3: Noise in Context. Since realistic projects usually consist of code with different functions, the code context is likely to contain noise code not entirely related to malicious purposes, which may have a potential impact on the effectiveness of the final induced perturbation. As such, we design the noise code, adding irrelevant noise to enhance the robustness of the induced perturbation.
Based on three logics above, we can construct the code contextual simulation for different malicious objectives. To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode. The task components are automatically generated with the following stages. First, the position code should contain a fundamental encryption situation: the encryption key “key”, the text to be encrypted “iv” and the first half of the encryption instruction “cipher = DES.new(key,”. These variables and functions are necessary for the encryption methods, e.g., “CBC”. To achieve the malicious objective, the complete code instruction should call the weaker encryption method, “ECB”, thus the target code is “DES.MODE_ECB)”. After obtaining the output tuple, ShadowCode then declares the Crypto series methods that enable the encryption code to run normally in the conditional code. Finally, ShadowCode constructs the code that is not related to the task and uses it as the noise code. The final code contextual simulation is constructed from multiple aspects for the same task to collaboratively ensure the effectiveness.
Figure 3: The example of code contextual simulation, where the attacker aims to induce the Code LLM to use encryption mode with low security (i.e., ECB encryption). The simulation consists of four code segments: conditional code, noise code, position code, and target code, to simulate the context environment most relevant to the malicious objective.
Loss Function of Perturbation Optimization. After obtaining the code contextual simulation, a loss function can be formulated for optimizing induced perturbations. To achieve the malicious objective, the optimization aims to identify the perturbation that maximizes the probability of realizing the simulated successful attack. Specifically, the simulated contextual code is first divided into input and output components. The input to the Code LLM comprises conditional code, noise code, and position code, simulating a code completion task that can be exploited to achieve the malicious objective. This entire input is referred to as task code. The output, termed target code, serves as the label for the perturbation optimization task. Subsequently, with $\bm{Y}_{T}=[t_{1},t_{2},\ldots t_{k}]$, the loss function is defined as:
$\mathcal{L}=$ $\sum_{i=1}^{|\bm{Y}_{T}|}\log(1-\mathbb{P}(t_{i}|\bm{T}_{s}\leftarrow\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (2)
where $\bm{T}_{s}$ represents the task code, the probability $\mathbb{P}$ is inferred and calculated using the surrogate model weights $\bm{\theta}$. This function aims to maximize the likelihood of generating the target code given the task code as input. When the Code LLM receives the perturbation, it is induced to generate the target code if the context is appropriate (i.e., similar to the simulation in code functionality), thereby achieving the attacker’s malicious objectives.
3.4 Perturbation Optimization
In this stage, we first introduce two enhancement modules for the optimization process: forward reasoning enhancement and keyword-based perturbation design. Then, we will describe the whole algorithm for induced perturbation optimization.
Forward Reasoning Enhancement. Directly employing the loss function defined in Equation (2) may lead to distortions during the autoregressive generation process. As shown in Figure 4, since token prediction results are based on previous sampling outcomes, the original loss function may deviate from actual conditions due to unexpected predictions of the first few tokens. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context (i.e., immediately after calling a demo function “TRG”). If the induced Code LLM fails to generate the first token “os”, the probabilities of generating subsequent tokens will diverge significantly.
Figure 4: Distortion during the autoregressive generation process. The attacker aims to insert a malicious deletion command after the position code, whereas the Code LLMs would typically generate the result printing command. During the perturbation optimization process, the generation probability of the first token in the target code, i.e., “os”, is not adequately optimized. Consequently, the Code LLM may not be successfully induced upon encountering the position code and generate an undesired token “print”. Subsequently, the failure to generate the first token compromises the successful generation of subsequent tokens.
To alleviate this problem, a forward reasoning enhancement module is introduced. Motivated by the aforementioned discussion that subsequent tokens are influenced by preceding tokens in an autoregressive generation manner, earlier tokens hold greater importance for the overall generation process of the target code. Therefore, the importance of the initial tokens for the loss value is increased by assigning them higher weights, thereby enabling the optimization objective to adapt to the autoregressive mechanism. This increase in weight can be achieved through adding an additional loss. Its specific formulation is shown as follows:
$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-\mathbb{P}(t_{i}|(\bm{T_{s}}\leftarrow\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (3)
$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
where $\mathcal{L}_{e}$ is the additional loss, $r$ is its coefficient, and $h$ is the number of enhanced tokens.
Figure 5: Examples of different induced perturbations. The attacker aims to induce the Code LLM to generate a data deletion command (“os.system("rm -rf /")”) immediately after a specified function call (“TRG”) via induced perturbation. Manually designed induced perturbations lack stealthiness due to their obvious malicious intent, a flaw that a simple greedy search-based replacement alone cannot rectify. Induced perturbations optimized with random initialization necessitate extensive token usage to ensure effectiveness. Our method combines optimized induced perturbations with a prefix (keyword), thereby achieving both effectiveness and stealthiness with limited tokens.
Keyword-based Perturbation Design. This module aims to significantly reduce the token count required for induced perturbations by offloading information into prefix tokens, which enhances attack stealth. Arguably, minimizing token usage and obscuring malicious characteristics are critical to maintaining attack stealth, thereby avoiding detection by victim users or RAG’s developers. Generally, perturbations optimized via greedy search typically exhibit less obvious malicious characteristics compared to manually designed ones. However, these perturbations struggle to accurately convey malicious instructions as effectively as manually designed perturbations, resulting in higher token usage. A seemingly straightforward solution, which involves optimizing based on manually designed perturbations by replacing obvious malicious characteristics with equivalent tokens, proves difficult in practice. As shown in Figure 5, the tokens constituting these obvious malicious characteristics (i.e., convey dangerous data deletion instructions) are the most critical and challenging to replace via greedy search.
To alleviate these limitations, we propose a simple yet effective enhancement method, i.e., keyword-based perturbation design. Intuitively, the purpose of the induced perturbation is to fulfill the attacker’s malicious objective. Therefore, the output tuple, which expresses malicious objective, will contain the necessary information for the induced perturbation. Motivated by this understanding, we leverage tokens from the output tuple to construct a prefix for the perturbation, termed a keyword, which is then injected into the victim’s code context alongside the perturbation. During keyword construction, we use the loss function in Equation (3) to greedily select the optimal token combination. In particular, we utilize only a very small number of tokens (typically one or two) for keyword construction to ensure no overt malicious intent is present in the injected content. As shown in Figure 5, the keyword “TR” exhibits no overt malicious intent but can enhance the perturbation’s effectiveness compared to directly optimized counterparts.
Optimization Algorithm. We hereby introduce the specific optimization process of ShadowCode for induced perturbations and their corresponding keywords. Specifically, we employ greedy gradient search to optimize an initialized perturbation $\bm{P}=[p_{1},p_{2},\ldots p_{l}]$, where $l$ denotes the token count. As shown in Algorithm 1, our ShadowCode method optimizes the perturbation token by token in each iteration (line 4). We utilize the loss function in Equation (3) for greedy gradient search (line 5), i.e., the optimization objective for the $i$-th token $p_{i}$ is:
$\begin{array}[]{c}p_{i}=\mathop{\arg\min}\limits_{p_{i}}\mathcal{L}.\end{array}$ (4)
To achieve this, ShadowCode propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $p_{i}$, it greedily searches for the suitable token $e^{\prime}_{i}$ within the Code LLM’s vocabulary set $V$ to replace $p_{i}$, as shown in the following equation:
$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-p_{i}\right)^{T}\nabla_{p_{i}}\mathcal{L},$ (5)
where $\nabla_{p_{i}}\mathcal{L}$ represents the gradient of the task loss. ShadowCode enhances this token replacement strategy by employing beam search, which considers the Top-$k$ token candidates (Line 6-7). ShadowCode then replaces the original token with each candidate to form a new $\bm{P^{\prime}}$ (Line 8-9), and computes a new loss value $\mathcal{L}^{\prime}$: if $\mathcal{L}^{\prime}$ is better than $\mathcal{L}$, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire iteration, ShadowCode considers it optimal and terminates the optimization process (Line 16).
Algorithm 1 Induced Perturbation Optimization
1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{Y}_{T}$, initial perturbation $\bm{P}$ with tokens $p_{1:l}$, beam search size $k$
2:  Vocab set $V$, weights $\bm{\theta}$ of the surrogate Code LLM
3:  repeat
4:     for $i\in[0,\ldots,l]$ do
5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{Y}_{T},\bm{P},\bm{\theta})$
6:        $\mathcal{G}=\nabla_{p_{i}}\mathcal{L}$
7:        $\mathcal{P}_{\text{candidates}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
8:        for $j\in[0,\ldots,k]$ do
9:           $\bm{P}^{\prime}:\text{replace }p_{i}\text{ to }\mathcal{P}_{\text{candidates}}^{j}$
10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{Y}_{T},\bm{P^{\prime}},\bm{\theta})$
11:           if $\mathcal{L}^{\prime}<\mathcal{L}$ then
12:              $\bm{P}=\bm{P}^{\prime}$
13:           end if
14:        end for
15:     end for
16:  until $\bm{P}$ does not change in the whole iteration
17:  Output: Optimized perturbation $\bm{P}$
Additionally, we use grid search to optimize the keyword. The perturbation and the keyword are optimized alternately, which can be expressed as follows:
$\bm{P}_{key}^{s+1}=\arg\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}^{s},\bm{\theta}),$ (6)
$\bm{P}^{s+1}=\arg\min_{\bm{P}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P},\bm{\theta}),$
where $s$ is the number of optimization steps and $\bm{P}_{key}$ represents the keyword. $\oplus$ means concatenating two token sequences. We select keyword tokens from the output tuple, i.e., $\forall p_{ki}$ in $\bm{P}_{key},p_{ki}\in\{\text{\emph{target code},},\text{\emph{position code}}\}$. For example, as shown in Figure 5, the keyword “TR” are tokens from the position code of the output tuple.
4 Experiments
4.1 Main Settings
TABLE I: The summary of objectives and datasets. We list the programming languages that each objective/dataset is applicable: C/C++ (C), Python (P), Java (J). The last row of the table shows the total number of test data we design. Some objectives apply only to C/C++, resulting in a total of 31 applicable cases.
Identifier/Name Applicable Languages
Objectives ST0, ST1, ST2 C/C++, Python, Java
CWE416, 122, 787, 843 C/C++
CWE20, 78, 502, 918, 22, 306 C/C++, Java
Datasets MBPP Python
Humaneval Python
Eval-Plus Python
CodeXGLUE Python,Java
Humaneval-x C/C++,Python,Java
Data Count: 14273 (C/P/J): 2132/7965/4176
TABLE II: The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in boldface.
[…]
  CodeGemma-2b No Attack 1.9% 17.3% 9.4% 22.0% 4.9% 12.6% 17.0% 12.2%
C-GCG [26] 40.6% 50.2% 32.1% 36.5% 49.4% 59.0% 44.2% 44.6%
IPI-M [18] 3.4% 30.8% 6.4% 33.3% 3.2% 19.7% 31.0% 18.3%
ShadowCode (Ours) 75.5% 93.4% 84.2% 74.4% 67.3% 94.1% 85.4% 82.0%
CodeGemma-7b No Attack 0.0% 12.2% 0.0% 18.9% 0.0% 15.6% 12.1% 8.4%
C-GCG [26] 43.5% 53.8% 55.4% 33.1% 20.4% 47.2% 50.6% 43.4%
IPI-M [18] 11.1% 29.6% 4.3% 46.8% 20.8% 37.8% 33.2% 26.2%
ShadowCode (Ours) 85.8% 97.9% 78.5% 81.6% 70.5% 94.4% 97.4% 86.6%
CodeGeeX2-6b No Attack 0.1% 13.4% 12.9% 19.0% 24.0% 21.9% 14.3% 15.0%
C-GCG [26] 42.9% 46.1% 44.9% 41.7% 32.5% 51.7% 33.7% 41.9%
IPI-M [18] 9.5% 11.4% 5.3% 17.9% 0.9% 22.0% 11.5% 11.2%
ShadowCode (Ours) 79.6% 95.0% 80.8% 83.3% 72.0% 97.9% 85.6% 84.9%
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [27] across 3 programming languages: Python, Java, and C/C++. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks [10], while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research [52, 15], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We construct cases using the method described in Section 3.3 based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. The details of these cases are discussed in Appendix A. In addition, we collect corresponding code completion datasets, including Humaneval [53], Humaneval-x [5], CodeXGLUE [54], MBPP [55], and Eval-Plus [56]. The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Some datasets do not cover Java or C/C++, as shown in Table I. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma [4] and CodeGeeX [5], for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [31] on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications [2] with more than 935k downloads (on VScode [28] marketplace). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale code-oriented models, such as CodeBERT [57] and CodeT5 [58], because large-scale code-oriented models perform much better than them in code completion tasks.
Baseline Selection. We compare our ShadowCode to GCG-based and manually designed IPI attacks. In particular, the vanilla GCG [26] is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work [18]. IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments. Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor. More details for baselines are provided in Appendix A.
Evaluation Metrics. In this paper, we use ASR to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness. For ASR, we consider an attack successful only when the generated code has equivalent functionality to the target code and appears in the correct position. In other words, the generated code must not only have malicious functionality but also adhere to the code compilation syntax rules. NBR and STF are new metrics introduced for our target-specific attack scenarios. NBR represents the rate of feasible cases. The more cases remaining after excluding the unusable ones with low ASR (below a threshold $r_{1}$), the greater variety of malicious objectives the attack can achieve; STF represents the rate of successful cases and their threat level. The more cases with sufficiently high ASRs (above a threshold $r_{2}$) and larger overall ASR value, the greater the upper limit of the attack’s threat. Their formulations are as follows:
NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$ (7)
STF $=\frac{\sum_{i\in\text{cases where ASR}>r_{2}}\text{ASR}_{i}}{\text{Total \# of cases}},$
where $r_{1}$ and $r_{2}$ are set to 20% and 80% respectively in our evaluation empirically. Further reasons of evaluation metrics selection are discussed in Appendix A.
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing induced perturbations is set to 100. For forward reasoning enhancement module, the number of tokens reinforced is set to $h=2$, as defined in Equation (3). For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate $r$ in Equation (3) is 0.4, the number of keyword is 2, and the perturbation length is 10. Tokenization to calculate token number is performed using Gemma’s tokenizer.
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hours for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality induced perturbations tailored to the respective model sizes across all the experiment threat cases.
4.2 Evaluation on Open-source Code LLMs
We first perform the attack evaluation on open-source Code LLMs. This includes three aspects: (1) we assess the attack effectiveness over 3 Code LLMs across all threat cases, demonstrating that ShadowCode is widely effective and more threatening than other attacks; (2) we evaluate attack flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
Performance on Effectiveness. We hereby evaluate whether ShadowCode can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For ShadowCode, we use its default settings: perturbation length of 10 and 2 keyword, while C-GCG uses a prefix length of 12. Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens. Table II provides the average ASR for four methods. The effectiveness of ShadowCode significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to ShadowCode, which may be attributed to its strong contextual learning ability. Comparing the three programming languages, Python is the most susceptible language to ShadowCode due to its concise syntax and the smallest number of tokens consumed; Besides, ShadowCode is relatively less effective on CodeXGLUE, which is mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of ShadowCode. In other words, when the contextual code becomes complex, the strength of ShadowCode (e.g., the length of induced perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section 4.3 and Section 4.5, increasing the length of ShadowCode or the Top-$k$ value can effectively enhance the attack performance of ShadowCode. Additionally, we further enrich our experimental results by exploring the attack effectiveness for each case in Appendix B.
TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs. Best results are marked in bold.
Metric Method C/C++ Python Java
  NBR No Attack 0.00 0.22 0.01
C-GCG 0.39 0.50 0.56
IPI-M 0.08 0.25 0.17
Ours 0.77 0.92 0.78
STF No Attack 0.00 0.11 0.03
C-GCG 0.26 0.47 0.39
IPI-M 0.09 0.19 0.07
Ours 0.76 0.87 0.74
Flexibility. Previous experiments indicate that ShadowCode exhibits the best average ASR performance across various models. However, in real-world scenarios, attackers may have more complex demands. Generally, these demands can be categorized into two situations. One is when attackers want to achieve diverse malicious objectives, which requires that cases with low ASR be minimized. The other is when attackers want to cause severe damage, which requires that cases with high ASR be maximized. We refer to the capability to meet these complex demands as ”flexibility.” NBR is used to evaluate attack flexibility for the first situation, and STF for the second.
As shown in Table III, ShadowCode has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, ShadowCode scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for ShadowCode, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, ShadowCode significantly outperforms the baselines.
(a) ASR/STF/NIC under Different NIT
(b) NIT/NIC under Different ASR
Figure 6: The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, NIT 0 (No Attack) and NIT 2 to 20 (with a sampling interval of 2).
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. ShadowCode can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect ShadowCode attacks. Therefore, we measure the stealthiness of ShadowCode by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. Before reaching 60% ASR, ShadowCode attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing ShadowCode’s performance does not require significantly increasing the length of the perturbation as with existing methods [18], every token can effectively improve the ASR.
Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR. This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information. The tokens used by ShadowCode do not have distinct characteristics in terms of character length but are determined by the attacker’s design of the token number. This means that filtering ShadowCode attacks based on character length is unreliable.
4.3 Evaluation on Commercial Code LLMs
In previous experiments, we demonstrate that ShadowCode reaches sufficient effectiveness, flexibility, and stealthiness on open-source models. However, some widely used models are commercial and therefore are closed-source in real-world scenarios. In these cases, attackers need to conduct transfer attacks, i.e., generating induced perturbations using open-source Code LLMs to attack closed-source ones. We hereby evaluate ShadowCode’s transferability by leveraging two representative real-world scenarios, thereby further verifying its practical threat.
Attack Scenarios on Real-world Applications. The scenarios with limited access to Code LLM information can be broadly divided into two categories: gray-box and black-box, with the corresponding explanations as follows:
  1. 1)
    Gray-Box Setting. The attacker can infer the open-source model used by the target application through interaction but cannot obtain the exact version of the model. For example, the scale of the model’s parameters, fine-tuning information, and version numbers are unknown to the attacker. We simulate this scenario using CodeGeeX2-6b and the CodeGeeX application. To the best of our knowledge, the CodeGeeX model has at least four different versions, and the CodeGeeX application does not disclose the specific model information it uses.
  2. 2)
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as Codex). We simulate this scenario by attacking GitHub Copilot (supported by Codex). To align with gray-box setting, we adopt CodeGeeX2-6b to generate induced perturbations.
TABLE IV: The evaluation of transferability. We test ShadowCode’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
Scenario Language ASR NBR STF
  Gray-box C/C++ 80.4% 0.92 0.75
Python 93.3% 1.00 0.93
Java 71.7% 0.78 0.63
Black-box C/C++ 67.7% 0.70 0.64
Python 90.4% 1.00 0.85
Java 74.4% 0.78 0.65
(a) Perturbation Injection
(b) Target Completion
Figure 7: The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a function. When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after calling “TRG”).
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section 4.2.
Results. As shown in Table IV, in the gray-box scenario, ShadowCode still demonstrates great attack performance. In Python and C/C++, ShadowCode achieves ASRs of 80.4% and 93.3%, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, ShadowCode also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our ShadowCode still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, ShadowCode achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, ShadowCode achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
In Figure 7, we present a representative case of ST2 in the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
4.4 Resistance to Potential Defenses
In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against ShadowCode.
Figure 8: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. T-SNE plot figures at the other layers have similar results.
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code that contains perturbations and normal code, thus rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach. This involves examining the differences between the benign code and the injected code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 8, we sample 1640 code segments and construct classification prompts. We detailly discuss the setting of this semantic analysis in Appendix B. The results show that benign code and injected code are almost completely indistinguishable.
Figure 9: The resistance to the removal of chars. We remove characters from the perturbation, selecting 5% for removal each time, and we calculate the average ASR for each programming language across applicable cases. As the influence of the perturbation fades, the ASR results converge to the same level as no attack baseline.
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect injected code segments (as mentioned above). To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 9, ShadowCode still maintains a sufficiently high ASR under a 50% removal rate. On the other hand, even if readability is completely sacrificed by extensively removing comments, ShadowCode can still be effective through non-comment portions of the code, as indicated in Table V. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since ShadowCode imposes strict limitations on the length of perturbations.
TABLE V: ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments.
[…]
Variable Assignment C/C++ 67.5% -8.0%
Python 90.1% -7.8%
Java 78.6% -5.6%
Output Content C/C++ 65.0% -10.5%
Python 83.0% -14.6%
Java 66.5% -9.0%
4.5 Ablation Study
In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for ShadowCode attacks. We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our ShadowCode.
(a) ASR Study
(b) STF Study
Figure 10: Ablation study of keyword number and Top-$k$. We compare ShadowCode’s ASR and STF with keyword length from 0 to 2. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
Effects of Top-$k$. This experiment validates the impact of Top value on the performance of ShadowCode. As shown in Figure 10, increasing the Top-$k$ value can significantly improve ShadowCode’s ASR and STF. Arguably, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the LLM’s vocab size. However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400).
(a) No FRE
(b) FRE
Figure 11: Ablation study of FRE. We compare the differences in ASR and losses both with and without using FRE module. 1-key1 and 1-key2 represent keeping one of the two selected keyword, respectively.
Effects of Keyword-based Designing. This experiment evaluates the effects of our optimization module, i.e., keyword-based designing. Specifically, we vary the number of keyword and generate the perturbations. As shown in Figure 10, limiting the number of keyword has a significant impact. The result of 1-keyword perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keyword perturbation. The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%. In contrast, 2-keyword perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95). The above results indicate that using keyword can effectively improve the performance of ShadowCode.
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using the same keyword. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 11, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on ASR than keyword-based design. These results verify the necessity of forward reasoning enhancement in our ShadowCode method.
5 Potential Limitations and Future Work
As the first work on automatically attacking Code LLMs via external prompt injection, we have to admit that ShadowCode still exhibits some potential limitations. Firstly, we currently select the keyword tokens solely from the output tuple via a simple grid search. Although this method can already select effective keywords, it may overlook potentially better solutions. Developing more efficient keyword optimization strategies is an important potential research direction. Secondly, we currently do not mandate that ShadowCode generate entirely natural-looking induced perturbations. Arguably, unreadable content (e.g., symbols, abbreviations, and custom function names) is common in code comments and frequently does not raise suspicion. Meanwhile, as shown in Figure 7, the perturbation generated by ShadowCode constitutes only a small portion of the tokens in the entire code file, rendering it difficult to detect. However, we also believe that natural-looking perturbations can enhance the stealthiness of attacks, and we will explore how to leverage LLMs as optimizers to achieve this in our future work. Thirdly, our work only focuses on the code completion scenario. We will investigate the performance of ShadowCode in other scenarios such as code summarization and code translation in future work.
6 Conclusion
In this paper, we introduced external prompt injection (EPI), a new attack paradigm against Code LLMs. EPI manipulates Code LLMs by injecting concise, non-functional induced perturbations into the victim’s code, which can spread through common dependencies. In particular, EPI does not necessitate control over the model’s training process and can achieve specific malicious objectives, posing more threatening risks over existing backdoor and adversarial attacks. Furthermore, we proposed a simple yet effective method (dubbed ShadowCode) for this attack paradigm, capable of automatically generating induced perturbations to achieve effective and stealth attacks. We evaluated ShadowCode across three popular programming languages and 31 representative cases. We showed that our ShadowCode can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all cases, using a 12-token non-functional perturbation. Our work alerted to a new serious security threat in the use of Code LLMs.
Acknowledgments
We would sincerely thank Prof. Kui Ren and Prof. Chun Chen from the Zhejiang University for their helpful comments and suggestions on an early draft of this paper.
Appendix A Additional Details
Indirect Prompt Injection against Code LLMs. In code domain, existing indirect prompt injection [18] (IPI) uses readable instructions to interfere with the model’s reasoning results, commanding the model to generate insecure code. In their show case, they place a large amount of disruptive commentary along with the obvious malicious code in the comment section, as depicted in Figure 12. We argue that this method exhibits very low feasibility. Direct human inspection or simple semantic analysis can easily detect these malicious comments.
Case Study. Here we show the details of our malicious objectives and threat cases. As shown in Table VI, we design 13 malicious objectives based on CWE top-10 KEV, existing research and famous destructive instruction. In the programming process, minor changes in the code can cause significant functional alterations. We roughly divide the causes of vulnerabilities in all malicious objectives into three categories: skipping critical process (CWE20, 78, 502, 918, 22, 306), wrong memory usage (CWE416, 122, 787, 843), and outputting dangerous snippets (ST0, 1, 2). All these vulnerabilities result in severe consequences.
Figure 13 presents three representative examples, encompassing diverse languages and objective types. The first example, CWE20 (Java), involves a program that attempts to fetch URL content and calculate its length. Secure programming practice dictates checking for null content before length calculation to prevent null pointer exceptions. However, post-attack, the Code LLM generates code that directly calculates the length after URL fetching, resulting in a null pointer exception if the URL is null. Null pointers can lead to severe consequences, including program crashes or data corruption. We categorize this situation as “skipping critical process.” The second example, CWE122 (C), concerns a program that attempts to pass content to a buffer (‘buf‘) via a command-line argument. Secure programming practice necessitates checking if the content exceeds ‘BUFSIZE‘ before assignment, thereby preventing overflow. Using ShadowCode, attackers can induce Code LLMs to perform direct assignments. Stack overflows access non-existent RAM space, leading to consequences such as system crashes and data leaks. We categorize this situation as “wrong memory usage.” This type of situation typically applies only to C/C++, as Java and Python offer automatic memory management. The third example, ST2 (Python), involves a program that attempts to call a function named TRG. Under normal circumstances, this call presents no issues, and the large language model generates suggestions based on contextual semantics or directly produces an ¡eos¿ token. Using ShadowCode, attackers can compel Code LLMs to generate the execution code ‘rm -rf /‘ immediately after calling the ‘TRG‘ function. If executed, this code results in the deletion of all data within the root directory. We categorize this situation as “outputting dangerous snippets.”
Figure 12: Greshake et al. [18] use a large amount of disruptive information and explicit malicious code snippets in the comments, which has a significant impact on the feasibility of the attack.
Baseline Details. IPI-M and C-GCG serve as crucial baselines in our effectiveness experiments. IPI-M mainly follows the existing IPI attack [18], employing human-readable instructions to compel the model to generate insecure code. However, due to the low stealth of the vanilla method, we simplify the instructions into concise, clear, template-based injection sentences, based on case categorizations. Specifically, for “skipping critical process,” we use “… directly after …” and “without any check”; for “wrong memory usage,” we use “It is important to … directly after …”; for “outputting dangerous snippets,”, we use “It is very important to promptly … after …”. Notably, the perturbation in IPI-M must be sufficiently explicit; for example, it needs to specify the code snippet you want to insert (e.g., “os.system("rm -rf")”) or present a clear assertion (e.g., “SSL3 is the most secure protocol”). Otherwise, Code LLMs are difficult to induce successfully. Arguably, the actual performance of IPI-M may be worse than our evaluation results shown in main experiments because of obvious malicious intent. C-GCG mainly follows the GCG method, employing an induced perturbation greedily optimized based on output tuple to achieve malicious objectives. Compared to ShadowCode, C-GCG does not use code simulation and two key modules: forward reasoning enhancement and keyword-based design.
Value and Metric Selection. We select the values of $r_{1}$ and $r_{2}$ based on the attack’s performance in the experiment. Specifically, in our experiments, 20% ASR is smaller than the ASR of no attack baseline in some situations, and 80% ASR outperforms the best performance of all the baselines. In addition, we find that only a few subcases (determined by the model, dataset, attack method, and case) have a success rate between 20% and 80%, accounting for 7.5%. Therefore, setting 20% and 80% as threshold of bad case and outstanding case respectively is appropriate.
As for metrics, we have already explained why metrics such as NBR and STF are used in the main experiments. Here we further explain some details related to them. First, NBR does not take the specific ASR value into account, as this metric is used to measure the diversity of attacks, rather than ensuring stable success rates. Second, STF targets the situation where attackers aim to cause as much damage as possible, and it varies with differences in ASR. In other words, even a difference between 100% ASR and 95% ASR can have a real impact on the overall threat level. Therefore, the calculation of STF considers the specific ASR value. In addition, we do not use common metrics for code generation tasks, such as BLEU or CodeBLEU [47], because they do not directly reflect the effectiveness of the attacks. Overall, we attempt to confirm the realistic meaning of each evaluation metric from different perspectives.
TABLE VI: Description of threat cases.
Case Description From
CWE416 Use after free. CWE 1st KEV
CWE122 Heap-based buffer overflow. CWE 2nd KEV
CWE787 Out-of-bounds write. CWE 3rd KEV
CWE20 Improper Input Validation CWE 4th KEV
CWE78 ‘OS Command Injection’ CWE 5th KEV
CWE502 Deserialization of untrusted data. CWE 6th KEV
CWE918 Server-side request forgery. CWE 7th KEV
CWE843 Access of resource using incompatible type. CWE 8th KEV
CWE22 ‘Path Traversal’ CWE 9th KEV
CWE306 Missing authentication for critical Function. CWE 10th KEV
ST0 Use ECB mode to encrypt. Schuster et al.[10]
ST1 Use SSL3 protocol Schuster et al.[10]
ST2 Executing “rm -rf / Famous destructive instruction
(a) Skipping critical process - CWE20
(b) Wrong memory usage - CWE122
(c) Outputting dangerous snippets - ST2
Figure 13: Three examples of ShadowCode attack cases. Code LLMs normally generate code segment without vulnerabilities, but they lose this ability after being attacked by ShadowCode. Attackers can modify the output content at the specified location through ShadowCode, and they can even generate malicious code snippets that have never appeared in the context, e.g., “os.system("rm -rf /")”.
Appendix B More Experiments
Figure 14: The ASR across different cases. The experimental results shown in this figure are the average values calculated from the effectiveness experiments involving 3 programming languages, 4 datasets, and methods excluding Clean Accuracy.
Effectiveness across Different Cases. In the main experiments, we have validated the outstanding attack performance of ShadowCode from three perspectives: model, dataset, and programming language. Here, we supplement the attack details for different attacks on a single case. Specifically, we use ASR to evaluate the performance of ShadowCode and the baselines across different cases. We present the specific average ASR for each case. As shown in Figure 14, ShadowCode achieves a certain level of effectiveness across all cases, with an overall average ASR of 83.4%. However, both C-GCG and IPI-M have cases where they are completely ineffective, especially for IPI-M. This demonstrates that directly using traditional GCG methods or manually designing perturbations is not effective in some scenarios. We believe this involves three reasons: (1) The baseline methods are not fully suitable for contexts with structured features, which is a fundamental characteristic of the code completion task; (2) The specific malicious objective include complex semantics such as target code, position code, and “disregarding security attributes,” which are sometimes not fully captured by the perturbations generated by the baseline methods; (3) The impact of keyword-based design and FRE; our ablation study proves that these two modules play a critical role.
According to the results, the most challenging to achieve is the straightforward case ST2. This case requires generating an obviously malicious command in a completely unrelated context. Another difficult case to implement is the CWE22 “Path Traversal,” which involves using unconventional paths like “../” to access files that are normally inaccessible. This case is relatively easier to achieve in Python, where the Code LLMs tend to suggest directly reading the path. However, it is more challenging to implement in Java and C/C++, as Code LLMs often only recommend checking the path rather than directly reading the file. CWE843 involves manipulating union members in a struct in C/C++. This represents a more complex malicious objective, where we attempt to manipulate one value within the union using an untrusted value. This is an unconventional operation that is difficult to describe using natural language, making it challenging for IPI-M to successfully attack. In contrast, methods such as ShadowCode and C-GCG, which convey malicious instructions through token selection, are able to achieve success. Arguably, ShadowCode has a clear advantage in attack performance for all the cases.
Semantic Analysis on Payload Code. Here we further discuss the settings of the semantic analysis experiment. We observe the classification features of code segment in the embedding layer of CodeGemma-2b. Specifically, we add the classification prompt keyword [CLS] to 820 original code snippets and 820 malicious payload code snippets totally. and then input them into the model to analyze their semantic features. We examine the t-SNE [59] distribution of the classification keyword [CLS] in the embedding layer, which is a widely used method for classification tasks. As shown in Figure 8, the features of the two types of text are almost completely intertwined, making it nearly impossible to distinguish between the injected code data and the original code data. In addition to observing the characteristics of layer 17 shown in Figure 8, we also examine the data features of layers 1 through 16. There is no significant difference observed in these layers compared to layer 17. There are several reasons for this situation. First, code data, including comments, tend to have lower fluency and coherence with high token complexity, which poses a significant challenge for semantic analysis tasks. Second, the perturbations used by ShadowCode occupy an extremely small number of tokens, exerting only a minor influence on the overall semantics of the injected data, making them hard to detect. Third, the semantic impact of ShadowCode differs from traditional injection attacks in that it does not require an extremely high intensity of information to counteract the model’s secure alignment; instead, it guides the model to generate vulnerabilities that could occur naturally, thus making detection through semantic analysis difficult. Our experiment results indicate that simple semantic analysis methods are not effective in defending against ShadowCode. This indicates that detecting code snippets containing malicious perturbations generated by ShadowCode through semantic analysis is not effective.